# {strip}[Uncaptioned image]

### v1

@@ 5.2. Performance Optimizations
We implemented a small set of optimizations to make target applications usable. Memory move and conversion. Our kernel implements memory move using ARMv8 assembly, as these are often bottlenecks in framebuffer rendering. In the user library, we implement pixel buffer conversions (e.g., YUV to RGB) using ARMv8 SIMD. These optimizations improve video playback framerate by nearly 3x (from under 10 FPS to around 30 FPS for 480p video). Range disk I/O. Our kernel’s buffer cache, inherited from xv6, supports only single-block operations, which suffices for xv6’s simple filesystem but bottlenecks FAT32’s multi-block access. This severely slows large file loads (e.g., DOOM assets, videos). To mitigate this, we bypass the buffer cache for FAT32 range accesses, directly interfacing with the SD card driver, reducing latency by 2–3x. (See Section 6 for details).
@@ 5.3. Support for C++ apps
Our userspace implements a minimal runtime conforming to ARM’s bare-metal ABI (BPAPI (armDocumentationx2013, ; osdevOSDevWiki, )). This includes crt0.c (wrapping main()), and crti.c/crtn.c (handling global constructors/destructors). With under 100 SLoC, the runtime supports our C++ applications (e.g. Blockchain) as well as enabling future C++ apps like ML or game engines.
@@ 5.4. Excluded OS Features
File crash-consistency were excluded due to the lack of compelling applications requiring crash consistency and the complexity of journaling mechanisms, which adds challenges to filesystem understanding. Network stack implementation, while feasible with our USB/Ethernet driver and available opensource TCP/IP stacks, was excluded as networking is not a course objective. Pthreads were omitted due to the lack of support in the recent newlib version and the complexity of popular implementations like glibc or musl (musllibcMuslLibc, ). Current parallel programs directly use syscalls for thread management. Signals and related syscalls were not implemented; this however limits our compatibility with tools like GDB. HDMI audio output was excluded as it requires a proprietary firmware blob, adding unnecessary complexity for our purposes.
@@ 5.5. Development and Deployment
VOS as handheld Leveraging the Pi3’s rich peripheral ecosystem, we pair it with Game HAT, a handheld expansion board (Figure 1(h)), which is more self-contained than other kits (e.g., RPi Personal Computer Kit, SmartiPi Touch). Game HAT’s buttons connect via GPIO and emit key events through /dev/events. With this setup, VOS can render to a built-in 3.5” display as well as external displays such as big TVs (Figure 1(i)) and classroom projectors (f). Section 6 evaluates battery life of VOS on Game HAT. QEMU Our toolchain and scripts run on Ubuntu 22.04. To support Windows and macOS users, we test the setup on WSL2 (Windows), VMware Player (Windows/x86), and VMware Fusion (Apple Silicon). We fixed a few QEMU bugs during development. Development Process We develop VOS in three main stages: (1) Implement all required features in a monolithic OS codebase; (2) Identify target apps and map them to essential OS features; decompose the OS into five prototypes accordingly; (3) Refine each prototype, define student experiments, and maintain them separately. We also introduce demo functionalities in the final prototype.
@@ 6. Evaluation
We test our OS prototypes with apps on Raspberry Pi 3 (Pi3 in short) and QEMU as listed in  Table 2.
@@ 6.2. Microbenchmarks
The building blocks of our OS exhibit low overheads, as shown in Figure 8: (1) The overhead of a syscall is around 3 $\mu$s. (2) One-way IPC between user processes takes approximately 21 $\mu$s, which includes scheduling, data copying, and context switch; (3) The FAT32 read/write throughput achieves several hundred KBs per second; the throughput is primarily constrained by our SD driver, which lacks DMA support and polls registers for read and write. (4) The entire device takes around 6 seconds from power-on to shell prompt; much of this time is spent by the firmware loading our kernel and our kernel initializing slow peripherals such as USB controllers.
@@ 6.3. App benchmarks
VOS delivers good performance for its target apps.
@@
Our benchmark configurations include: (1) DOOM which uses direct rendering without a window manager; (2) Video player which decodes both video and audio streams and uses direct rendering; video files are preloaded into memory before decoding. (3) Mario in three variants. (i) mario-noinput (Prototype 3) which runs one thread and uses direct rendering, without handling any input events; (ii) mario-proc (Prototype 4) which runs multiple processes to handle input events; it uses direct rendering; (iii) mario-sdl (Prototype 5) which runs multiple threads to handle the events; it renders indirectly via a window manager.
@@
(Figure 11) In these experiments, we make DOOM and mario variants render as fast as possible without locking to a fixed FPS; we make video playback target a video’s native framerate. FPS is measured after a 20-second warm-up period. As illustrated in  Figure 11, both DOOM and mario on VOS achieve decent throughput of 62 – 115 FPS. In video playback, our system attains 27 FPS for 480p video, close to its native framerate (30 FPS), while considerably lower for 720p video is (12FPS). Among three mario variants, mario-noinput has similar FPS to mario-sdl, which are both lower than mario-proc; the differences reflects the impact of our OS designs, which will be analyzed in latency breakdown below. All benchmarks see higher throughts (from 13% to 1.5x) when VOS runs on QEMU atop a modern x86 machine.
@@
We further compare our throughputs to Ubuntu/Linux and FreeBSD (configurations in Table 3). They run on the same hardware (Pi3), use the same compiler and optimization levels (gcc-11 with O2), and build the same app sources. For minimum overhead, we run Linux and FreeBSD on X servers without a window manager. The results show that our throughputs are comparable to Linux/FreeBSD, ranging from 0.8x to 1.9x, or in the difference of -15–30 FPS. Note that we do not claim that our system is competitive against production OSes, as the latter is capable of hardware-accelerated rendering which our apps do not utilize; and that the latter incurs the overhead of a generic X server which our system eschews.
@@
Rendering latency
Figure 12(a) reveals the sources of rendering overhead. Overall, rendering delays are dominated by app execution: primarily the application logic such as game engine and video/audio decoding and drawing (e.g. buffer and pixel manipulation); by contrast, the kernel incurs low overhead, e.g. writes to mmap’d framebuffer or device files. The latency also explains the FPS difference when running the three mario variants on our OS (Figure 11): mario-sdl shows higher latency in its app logic than other two variants, likely because inclusion of a full-fledged C library (newlib) adds overhead to its game execution.
@@
Overall, such end-to-end input latency corresponds to as low as 1–2 game frames. A detailed breakdown in Figure 12(b) shows that: (1) the majority of the latency comes from the app logic: both DOOM and Mario poll for input events in their main event loops, in which polling intervals (5 to 13 ms) contribute noticeable latency. (2) Among the remaining components, the extra event indirection (from window manager for mario-sdl, and the IPC between main/input tasks for mario-proc) contribute significantly, which is the cost for OS modularity. (3) Notably, our kernel’s pipe(), a simplistic design ported borrowed from xv6, becomes a bottleneck even for passing a keyboard event (fewer than 10 bytes). DOOM, without such event indirection, shows the lowest latency.
@@
Our system scales well with the number of CPU cores, on a multi-programmed workload (eight simultaneous Mario instances, screenshot in Figure 1(l)) as well as on a multi-threaded workload (Blockchain miner). As shown in  Figure 11, as the number of cores increases, the throughput sees proportional growth. We also verify that the utilization of all CPU cores are constantly over 95%.
@@
VOS has low memory usage. When running target apps individually (e.g., Mario, DOOM, and video player), we measured total OS memory usage between 21 MB and 42 MB — about 2% to 4% of the Pi3’s 1 GB of DRAM.
@@ 6.4. Power efficiency and battery life
Figure 14. A pedagogical survey on VOS, validating our design principles P1–P4 (§1). Responses on a scale of 1 (strong disagreement) to 5 (strong agreement). N=48.
@@ 6.5. Pedagogical user study
In Spring 2025, we offered an OS course centered around VOS for the first time. Toward the end of the semester, we conducted a survey and received 48 responses out of 59 enrolled students, who were a mix of computer science and engineering majors.
Figure 14 summarizes students’ ratings on how VOS’s four principles (P1–P4, §1) directly supported their learning. (P1) Most students found the interactive apps to be a strong motivator. Many appreciated the opportunity to showcase their apps to others. (P2) Although students could choose between working on real hardware or using emulators, a majority (64%) opted to experiment with real devices and were motivated. Some students noted challenges in setup and debugging. (P3) Incremental prototyping was effective, with early prototypes serving as clear stepping stones to later ones. (P4) Students reported a clear understanding of the relationship between apps and their dependencies, as well as the connections between different prototypes. Overall, the survey validates the pedagogical value of VOS.
@@ 7. Conclusions
VOS showcases that building a modern, usable OS can be both demonstration-ready and educational. By tying modern OS features to real apps through inverse engineering, we keep the OS construction a motivated journey, with a clear purpose behind system design. The result is a small, efficient kernel that supports rich apps and runs on real hardware. VOS makes OS construction more approachable for a wider audience, without losing depth or realism.

### v2

@@ 5.2. Performance Optimizations
We implemented a small set of optimizations to make target applications usable. Our kernel implements memory move using ARMv8 assembly, as these are often bottlenecks in framebuffer rendering. In the user library, we implement pixel buffer conversions (e.g., YUV to RGB) using ARMv8 SIMD. These optimizations improve video playback framerate by nearly 3x (from under 10 FPS to around 30 FPS for 480p video). Our kernel’s buffer cache, inherited from xv6, supports only single-block operations, which suffices for xv6’s simple filesystem but bottlenecks FAT32’s multi-block access. This severely slows large file loads (e.g., DOOM assets, videos). To mitigate this, we bypass the buffer cache for FAT32 range accesses, directly interfacing with the SD card driver, reducing latency by 2–3x. (See Section 7 for details).
@@ 5.3. Support for C++ apps
Our userspace implements a minimal runtime conforming to ARM’s baremetal ABI (BPAPI  (armDocumentationx2013, ; osdevOSDevWiki, )). This includes crt0.c (wrapping main()), and crti.c/crtn.c (handling global constructors/destructors). With under 100 SLoC, the runtime supports our C++ applications (e.g. Blockchain) as well as enabling future C++ apps like ML or game engines.
@@ 5.4. Excluded OS Features
File crash-consistency was excluded due to the lack of compelling applications requiring crash consistency and the complexity of journaling mechanisms, which adds challenges to filesystem understanding. Network stack implementation, while feasible with our USB/Ethernet driver and available opensource TCP/IP stacks, was excluded as networking is not a course objective. Pthreads were omitted due to the lack of support in the recent newlib version and the complexity of popular implementations like glibc or musl  (musllibcMuslLibc, ). Current parallel programs directly use syscalls for thread management. Signals and related syscalls were not implemented; this however limits our compatibility with tools like GDB. HDMI audio output was excluded as it requires a proprietary firmware blob, adding unnecessary complexity for our purposes.
@@ 5.5. Development and Deployment
Leveraging the Pi3’s rich peripheral ecosystem, we pair it with Game HAT, a handheld expansion board (Figure LABEL:fig:collage(h)), which is more self-contained than other kits (e.g., RPi Personal Computer Kit, SmartiPi Touch). Game HAT’s buttons connect via GPIO and emit key events through /dev/events. With this setup, Proto can render to a built-in 3.5” display as well as external displays such as big TVs (Figure LABEL:fig:collage(i)) and classroom projectors (f). Section 7 evaluates battery life of Proto on Game HAT.
Our toolchain and scripts run on Ubuntu 22.04. To support Windows and macOS users, we test the setup on WSL2 (Windows), VMware Player (Windows/x86), and VMware Fusion (Apple Silicon). We fixed a few QEMU bugs during development.
We develop Proto in three main stages: (1) Implement all required features in a monolithic OS codebase; (2) Identify target apps and map them to essential OS features; decompose the OS into five prototypes accordingly; (3) Refine each prototype, define student experiments, and maintain them separately. We also introduce demo functionalities in the final prototype.
6. Student assignments
Lab #Tasks #Files SLoC #Videos
Lab1 13 10 $\sim$100 9
Lab2 10 10 $\sim$100 9
Lab3 7 18 $\sim$150 6
Lab4 7 21 $\sim$300 7
Lab5 6 28 $\sim$300 6
Table 2. Student workload for labs, showing numbers of tasks, numbers of source files to modify, lines of code to write, and numbers of required video evidences for assessment. See Figure 15 for detailed lab tasks.
6.1. Lab design
Throughout the semester, students complete five labs, each corresponding to one of the five prototypes described in Section 1. For each lab, students submit one assignment. We provide starter code for every lab, designed as a clean slate so that performance in an earlier lab does not affect grading in later labs.
Task structure
Assuming students have at most 150 hours for assignments (§2), we design labs to prioritize breadth of OS concepts over depth in any single component. Our goal is for students to work with many parts of an OS, with each part requiring relatively small (yet still crucial) tweaks or fixes. Each lab consists of multiple tasks, where each task exercises specific OS concepts, as shown in Figure 15. Tasks often have dependencies, and completing a sequence of dependent tasks unlocks interesting application features along the way. Students are not required to complete all tasks to earn full credit for a lab; they may choose their own path through the task graph based on the features they aim to implement. Some students find this structure analogous to video games, with a main storyline complemented by optional side quests.
Task workload
We, as instructors, create the starter code by removing key pieces from a working prototype. Each task is intentionally small, typically requiring students to understand 1–2 source files, modify several or more code locations, and write tens of lines of code in total. For example, one task asks students to complete the user-level memory mapping code so that mario can render graphics. Table 2 summarizes the typical student workload. This contrasts with many instructional OS assignments that require implementing an entire standalone subsystem (e.g., a filesystem).
Teamwork
Labs 1–3 are completed individually, while Labs 4 and 5 expect teamwork. The rationales are: (1) Labs 4 and 5 involve rich userspace and GUI applications, which may exceed the time budget of some students working alone; (2) many tasks in Labs 4 and 5 have inherent parallelism (e.g., separate kernel and user-level components needed by a window manager), allowing teammates to work concurrently and later integrate their work.
6.2. Student deliverables & assessments
We find it challenging to automatically test a custom OS running in QEMU or on real hardware. Thankfully, our labs are centered on media-rich applications; we thus require video evidence for critical tasks (marked with bold borders in Figure 15), making students responsible for demonstrating that their OS enables the required features (e.g., rendering four rotating donuts at different rates). Video submissions must meet strict criteria: minimum duration (e.g., 30 seconds), minimum resolution (e.g., 1080p), continuous display of the student’s university ID, and no post-editing. If video evidence is not provided, instructors will manually assess the student’s source code.
@@ 7. Evaluation
We test our OS prototypes with apps on Raspberry Pi 3 (Pi3 in short) and QEMU as listed in  Table 3. The evaluation shows our system (1) is capable of running feature-rich apps, (2) is competitive against production OSes on target apps, not only just microbenchmarks, (3) is sufficiently efficient for demonstration and use.
@@ 7.2. Microbenchmarks
[…]
@@ 7.3. App benchmarks
Proto delivers good performance for its target apps.
@@
Our benchmark configurations include: (1) DOOM, which uses direct rendering without a window manager. (2) VideoPlayer, which decodes both video and audio streams and uses direct rendering; video files are preloaded into memory before decoding. (3) mario in three variants. (i) mario-noinput (Prototype 3), which runs one thread and uses direct rendering, without handling any input events. (ii) mario-proc (Prototype 4), which runs multiple processes to handle input events; it uses direct rendering. (iii) mario-sdl (Prototype 5), which runs multiple threads to handle the events; it renders indirectly via a window manager.
@@
(Figure 11) In these experiments, we make DOOM and mario variants render as fast as possible without locking to a fixed FPS; we make video playback target a video’s native framerate. FPS is measured after a 20-second warm-up period. As illustrated in  Figure 11, both DOOM and mario on Proto achieve decent throughput of 62–115 FPS. In video playback, our system attains 27 FPS for 480p video, close to its native framerate (30 FPS), while considerably lower for 720p video (12FPS). Among three mario variants, mario-noinput has similar FPS to mario-sdl, which are both lower than mario-proc; the differences reflect the impact of our OS designs, which will be analyzed in latency breakdown below. All benchmarks see higher throughputs (from 13% to 1.5x) when Proto runs on QEMU atop a modern x86 machine.
@@
We further compare our throughputs to Ubuntu/Linux and FreeBSD (configurations in Table 4). They run on the same hardware (Pi3), use the same compiler and optimization levels (gcc-11 with O2), and build the same app sources. For minimum overhead, we run Linux and FreeBSD on X servers without a window manager. The results show that our throughputs are comparable to Linux/FreeBSD, ranging from 0.8x to 1.9x, or differing by 15-15–30 FPS. Note that we do not claim that our system is competitive against production OSes, as the latter is capable of hardware-accelerated rendering which our apps do not utilize; and that the latter incurs the overhead of a generic X server which our system eschews.
Rendering latency
Figure 12(a) reveals the sources of rendering overhead. Overall, rendering delays are dominated by app execution: primarily the application logic such as game engine and video/audio decoding and drawing (e.g. buffer and pixel manipulation); by contrast, the kernel incurs low overhead, e.g. writes to mmap’d framebuffer or device files. The latency also explains the FPS difference when running the three mario variants on our OS (Figure 11): mario-sdl shows higher latency in its app logic than the other two variants, likely because inclusion of a full-fledged C library (newlib) adds overhead to its game execution.
@@
Overall, such end-to-end input latency corresponds to as low as 1–2 game frames. A detailed breakdown in Figure 12(b) shows that: (1) the majority of the latency comes from the app logic: both DOOM and mario poll for input events in their main event loops, in which polling intervals (5 to 13 ms) contribute noticeable latency. (2) Among the remaining components, the extra event indirection (from window manager for mario-sdl, and the IPC between main/input tasks for mario-proc) contribute significantly, which is the cost for OS modularity. (3) Notably, our kernel’s pipe(), a simplistic design ported from xv6, becomes a bottleneck even for passing a keyboard event (fewer than 10 bytes). DOOM, without such event indirection, shows the lowest latency.
@@
Our system scales well with the number of CPU cores, on a multi-programmed workload (eight simultaneous mario instances, screenshot in Figure LABEL:fig:collage(l)) as well as on a multi-threaded workload (Blockchain miner). As shown in  Figure 11, as the number of cores increases, the throughput sees proportional growth. We also verify that the utilization of all CPU cores is constantly over 95%.
@@
Proto has low memory usage. When running target apps individually (e.g., mario, DOOM, and VideoPlayer), we measured total OS memory usage between 21 MB and 42 MB — about 2% to 4% of the Pi3’s 1 GB of DRAM.
@@ 7.4. Power efficiency and battery life
Figure 14. A pedagogical survey on Proto, validating our design principles P1–P4 (§1). Responses on a scale of 1 (strong disagreement) to 5 (strong agreement). N=48.
@@ 7.5. Pedagogical user study
In Spring 2025, we offered an OS course centered around Proto for the first time. Toward the end of the semester, we conducted a survey and received 48 responses out of 59 enrolled students, who were a mix of computer science and engineering majors.
Quantitative results
Figure 14 summarizes students’ ratings on how Proto’s four principles (P1–P4, §1) directly supported their learning. (P1) Most students found the interactive apps to be a strong motivator. Many appreciated the opportunity to showcase their apps to others. (P2) Although students could choose between working on real hardware or using emulators, a majority (64%) opted to experiment with real devices and were motivated. Some students noted challenges in setup and debugging. (P3) Incremental prototyping was effective, with early prototypes serving as clear stepping stones to later ones. (P4) Students reported a clear understanding of the relationship between apps and their dependencies, as well as the connections between different prototypes. Overall, the survey validates the pedagogical value of Proto.
Qualitative results
Students provided open-ended feedback on Proto and the course. They praised the following aspects: (1) strong application motivations (“Five OS prototypes… culminating in DOOM… kept me motivated”); (2) incremental complexity across prototypes (“The progressive nature of the labs… created a satisfying learning arc”); (3) diverse learning paths (“Choose-your-own-adventure format made labs digestible”); and (4) tight lecture–lab integration (“One without the other would’ve left me confused”). They also reported the following challenges: (1) difficulty of developing on real hardware (“RPI3 impractical to develop while traveling…”); (2) a desire for deeper exploration of certain topics (“Labs went for breadth over depth… wished side quests added depth”); (3) workload pressure (“Painful at times… worth it, but labs took significant time”); and (4) a need for more scaffolding in early labs (“Jump from basics to implementation was steep”). We will address this feedback in future course iterations.
Discussion
Looking ahead, we plan to continue refining the prototypes and incorporating feedback from students. To make the learning curve more approachable, we aim to enhance the documentation for each prototype, offering clearer background, motivation, and references to support student understanding. To broaden the learning experience, we plan to develop additional applications that integrate diverse kernel features, expand the set of lab branches so students can pursue different paths, and provide a curated list of project options for students to explore in their final work.
@@ 8. Conclusions
Proto showcases that building a modern, usable OS can be both demonstration-ready and educational. By tying modern OS features to real apps, we keep the OS construction a motivated journey, with a clear purpose behind system design. The result is a small, efficient kernel that supports rich apps and runs on real hardware. Proto makes OS construction more approachable for a wider audience, without losing depth or realism.
ACKNOWLEDGMENT
The authors were supported in part by NSF awards #2128725, #1919197, #2106893, and #2426353. The authors thank the anonymous reviewers and the shepherd, Eddie Kohler, for their insightful feedback.
Appendix A Appendix
We include the task graphs for labs 1–5 below.
Figure 15. Task graphs for all labs. Each box represents a task, with the corresponding OS concepts indicated in parentheses. Boxes with borders denote tasks for which students must submit video evidence for assessment.